# Combinatorial summation of Feynman diagrams: Equation of state of the $2D$ $SU(N)$ Hubbard model

### v3

@@ II.3 Approach 2: constructing connected diagrams from the start
There is much freedom in how the graph summing a particular series is designed, and the following general principles can aid its efficiency: (i) allowing unwanted or unphysical sequences $\mathcal{C}$ might be useful if they cancel in the final sum, and (ii) walks can traverse the cycle covers $\mathcal{C}$ and be grouped in arbitrarily order, provided all the required sequences are generated in the end. Principle (i) was key for computing the determinant, but it has another use here: we can formally allow on-site interactions between same-spin fermions in the Hamiltonian (1) since the resulting diagrams cancel. Instead of the topology-dependent factor $M$, the diagrammatic rules Abrikosov et al. (2012) for fully spin-symmetric interactions prescribe that each fermionic loop is multiplied by the number of spins, implying a multiplication of each node that closes a cycle merely by $N$. Although having to construct diagrams that cancel is a hindrance at lower orders, the simpler diagrammatic rules allow for a more efficient scaling at $n\gtrsim 5-6$.
@@
Our recipe for organising the walks that constitute the graph has so far been borrowed from the determinant, forcing us to keep track in $\mathcal{R}$ of how different cycles are connected. This is not necessary if we reorganise the walks to generate only connected diagrams from the start. Since for generic Hamiltonians we cannot rely on the cancellation of terms with repeated elements, we at least must keep track of the elements visited up to the current step $l$, =$\mathcal{R}=\{e_{1},e_{2},\ldots e_{l}\}$, and ban adding $e$ to $\mathcal{C}$ if $e\in\mathcal{R}$. Demoting the role of $h$ in the node label $[l,h,e]\otimes\{e_{1},e_{2},\ldots e_{l-1},e\}$ to being merely the first element in the current cycle, we can generate only connected diagrams if each new cycle starts with the element that is paired by an interaction to one of the already visited ones $e_{i}\in\{e_{1},e_{2},\ldots e_{l-1},e\}$, e.g. the smallest in $\mathcal{R}$ that is not already paired, for uniqueness. It is easy to see that the number of floating point operations in this graph is only exponential, $\mathcal{O}(n^{3})4^{n}$, and that the information about visited elements carried in $\mathcal{R}$ is minimal for this order of traversing the sequences $\mathcal{C}$, i.e. the graph cannot be pruned any further. The computational cost of this algorithm, labelled CoS-2, is shown for the $SU(N)$ case in Fig. 2b.
@@
In systems where there is no non-trivial factor associated with each fermionic loop, as, e.g., in the $SU(2)$ Hubbard model, or for $N=1$, cancellations between cycle covers with repeated elements can still be utilised to reduce the cost further to $\mathcal{O}(n^{3})3^{n}$. To this end, $\mathcal{R}$ only needs to store the list of interactions that a visited element belongs to, and whether only one vertex of the interaction or both have been visited, i.e. $3$ possibilities for each interaction. Since there is no record of which of the two vertices of an interaction has been visited, both options for the element that starts a new cycle need to be allowed, with the cycle cover that ends up repeating the vertex cancelling out, as in the case of the determinant. The complexity of this algorithm is plotted for the $SU(2)$ case in Fig. 2a.
@@
Finally, sums of skeleton (bold-line) diagrams in arbitrary channels can be straightforwardly generated in our approach. For instance, the computational cost of producing an expansion in terms of the full (interacting) Green’s function $G$ and screened interaction $W$  Hedin (1965) by a simple extension of the CoS-2 algorithm is plotted in Fig. 2b as CoS-GW. The challenge of restricting the series to irreducible diagrams in both channels is met here by supplementing the nodes in the CoS-2 graph with a record $\mathcal{R}$ that keeps track of connectivity when a propagator or interaction line is cut, similarly to the CoS-1 approach of Sec. II.2. Curiously, there is no notable cost increase relative to the CoS-1 for connected diagrams. The versatility of the CoS platform could enable more efficient algorithms for skeleton series in the future.
@@ II.5 Nonsymmetrised sum
Like the determinant, in its general form the CoS approach sums not just all diagram topologies of a given order $n$ but their $2^{n}n!$ realisations with all permutations of the interaction lines and end points. An advantage of the explicit summation is that this symmetrisation can be suppressed if useful with a reduction of the computational cost. Specifically, since the Hubbard interaction is local in space and time, the symmetrisation over $2^{n}$ exchanges of the interaction end points does not affect the variance of the Monte Carlo sampling. In this case, ordering the arguments in each $V(\alpha,\beta)$ leads to a tangible speed-up of the DiagMC calculation. Indeed, by requiring that a node $[l,h,\bar{e}]\otimes\mathcal{R}$ is added to the graph only if $e$ (the vertex connected to $\bar{e}$ by an interaction) is already present in $\mathcal{R}$, the number of operations in the CoS-2 approach drops from $\mathcal{O}(n^{3}4^{n})$ to $\mathcal{O}(n^{3}3^{n})$. The actual computational cost shown in Fig. 3 (CoS-2, ordered interaction ends) follows this scaling.
@@
Figure 3: Number of floating-point operations (FLOP) required to evaluate the sum of the integrands of all Feynman diagrams of order $n$ by the CoS-2 algorithm without inherent symmetrisation over vertex permutations: for connected diagrams in the $SU(N)$ case with ordered end points of interactions (CoS-2, ordered interaction ends) and fully ordered vertices (CoS-2, all vertices ordered). Shown as CoS-GW, all vertices ordered is the computational cost of summing the nonsymmetrised skeleton (bold-line) series in terms of the renormalised Green’s function $G$ and screened interaction $W$.
III Results
We compute the average particle number per lattice site $\langle n\rangle$ as a function of the chemical potential $\mu$, expressed in our approach as an expansion in the powers of the Hubbard coupling $U$, $\langle n\rangle(T,\mu,U)=\sum_{m=0}^{\infty}a_{m}(T,\mu)U^{m}$, in the thermodynamic limit. The series coefficients $a_{m}$ are obtained by the Monte Carlo integration of all connected diagrams of order $m$ over the positions of the $2m$ vertices in space-imaginary-time. Thus, $a_{m}$ are known numerically exactly with statistical error bars, while the only source of systematic error is the truncation of the series at order $n$. Although the series turns out to be divergent in all regimes of interest, being able to evaluate $a_{m}$ up to $n=8$ with $<5\%$ statistical error (and fractions of percent at lower orders) enables an accurate reconstruction of the result with controlled precision.
@@ III Results
The recent study by Pasqualetti et al. Pasqualetti et al. (2024) has revealed a perfect agreement between the DQMC calculations and experimental measurements of the EoS of the $2D$ $SU(N)$ Hubbard model down to $T/t=0.3$ and a coupling value up to $U/t=2.3$ for $N=6$. Fig. 4a shows the partial sum for the density at these parameters and $\mu/t=0.575$ as a function of the truncation order $n$. The series is seen to wildly diverge, but its analytic structure is rather simple, with the dominating singularity at $U_{c}/t\approx-0.9(1)$, which allows us to reconstruct the result following the approach developed in Ref. Šimkovic and Kozik (2019). Specifically, we employ the Dlog-Padé Baker Jr (1961) technique, taking into account the statistical error bars of $a_{m}$ and making sure that the systematic error of the evaluation—detected by the variation of the answer with free parameters of Dlog-Padé—is negligible. The result for the series in Fig. 4 is $\langle n\rangle=1.335(5)$. As a benchmark, we proceed to obtain the $\langle n\rangle(\mu)$ curve at $T/t=0.3$, $U/t=2.3$, plotted in Fig. 5, and find it to be in perfect agreement  [^2] The EoS reported in Ref. Pasqualetti et al. (2024) is averaged over small intervals of the chemical potential due to the experimental resolution, but we have verified the agreement within combined error bars for the prime numerical data, kindly provided by the authors. with that computed and measured in Ref. Pasqualetti et al. (2024).
[…]
Figure 5: Equation of state for the $2D$ $SU(N)$ Hubbard model with $N=6$ for $T/t=0.3,0.15$ and $U/t=2.3,4,8$.
Ibarra-García-Padilla et al. Ibarra-García-Padilla et al. (2021) demonstrate that the sign problem in DQMC rapidly intensifies with lowering $T$ and increasing $U$ and $N$ at the considered densities, as long as the system remains compressible. To explore the more challenging regime, the EoS was obtained by DiagMC at a lower temperature $T/t=0.15$ (see Figure 5), for which the $\langle n\rangle(\mu)$ curve is below that for $T/t=0.3$, indicating that the system is in the metallic state at $U/t=2.3$ and in this range of $\mu$ Lenihan et al. (2021). We further evaluate the series at larger values of $U$ up to $U=8$, where a faint shoulder around $\langle n\rangle=1$ is seen to emerge. This is consistent with the development of a (pseudo-)gapped state. At these couplings, the systematic error of resummation beyond the convergence radius becomes comparable to the propagated statistical error, and the combined error bar (a sum of the two errors) shown in Fig. 5 grows substantially. Nonetheless, the analytic structure of the series appears free from singularities near a positive real $U$, such as those in the $SU(2)$ Hubbard model at similar parameters Kim et al. (2020). There, the growth of the antiferromagnetic (AFM) correlation length beyond $\sim 10$ lattice sites was shown to be responsible for a near-critical behaviour of the diagrammatic expansions at these temperatures and $\langle n\rangle=1$ already at $U/t\sim 3$. Also in contrast to the $SU(3)$ case, where an insulating AFM ground state at $\langle n\rangle=1$ emerges already at $U/t\approx 5.5$ Feng et al. (2023) and strong AFM correlations (with a transformation upon heating) are observed up to $T/t\sim 0.5$ at $U/t=8$ Ibarra-García-Padilla et al. (2023), for $N=6$ AFM correlations appear weak down to $T/t=0.15$ at this coupling. Thus, there is no fundamental difficulty to reduce the errors bars and access larger $U$ values at the expense of a polynomially-longer calculation.
IV Discussion
The example application to the EoS of the $2D$ $SU(6)$ Hubbard model provides controlled benchmarks for ongoing theoretical and experimental studies, aimed at accessing lower temperatures and novel quantum many-body states. As a byproduct of a diagrammatic calculation, the analytic structure of the series offers additional insights in the physics of the system. The results suggest a phase transition in the attractive $SU(N)$ Hubbard model at a coupling strength as low as $U_{c}/t\sim-1$ up to temperatures $T/t\lesssim 0.5$, and absence of strong AFM correlations in the repulsive case at the considered temperatures and interaction strengths, at which the $SU(2)$imkovic et al. (2020); Kim et al. (2020) and $SU(3)$ Ibarra-García-Padilla et al. (2023); Feng et al. (2023) Hubbard models are already in the (quasi-)AFM state. In the $SU(2)$ case, the formulation in the thermodynamic limit enabled DiagMC to attain controlled accuracy in the regime where correlations are intrinsically long-range and are difficult to capture reliably by finite-size methods even in the absence of the sign problem Šimkovic et al. (2020); Kim et al. (2020). Such regimes of the $SU(N)$ model is where the developed technique can prove particularly useful. The possibility of a direct calculation of entropy in the DiagMC approach Lenihan et al. (2021) could be instrumental for thermometry in experiments with ultracold atoms that are currently testing the limits of state-of-the-art theoretical methods.
@@ Acknowledgements.
The author is grateful to Kaden Hazzard for illuminating and stimulating discussions, to Eduardo Ibarra-García-Padilla, Sohail Dasgupta, Kaden Hazzard, and Richard Scalettar for sharing their DQMC data, and to Sohail Dasgupta, Simon Fölling, Kaden Hazzard, Eduardo Ibarra-García-Padilla, Giulio Pasqualetti, and Richard Scalettar for a fruitful exchange of results and ideas. This work was supported by EPSRC through Grant No. EP/X01245X/1.

### v4

@@ II.3 Approach 2: constructing connected diagrams from the start
There is much freedom in how the graph summing a particular series is designed, and the following general principles can aid its efficiency: (i) allowing unwanted or unphysical sequences $\mathcal{C}$ might be useful if they cancel in the final sum, and (ii) walks can traverse the cycle covers $\mathcal{C}$ and be grouped in arbitrary order, provided all the required sequences are generated in the end. Principle (i) was key for computing the determinant, but it has another use here: we can formally allow on-site interactions between same-spin fermions in the Hamiltonian (1) since the resulting diagrams cancel. Instead of the topology-dependent factor $M$, the diagrammatic rules Abrikosov et al. (2012) for fully spin-symmetric interactions prescribe that each fermionic loop is multiplied by the number of spins, implying a multiplication of each node that closes a cycle merely by $N$. Although having to construct diagrams that cancel is a hindrance at lower orders, the simpler diagrammatic rules allow for a more efficient scaling at $n\gtrsim 5-6$.
@@
Our recipe for organising the walks that constitute the graph has so far been borrowed from the determinant, forcing us to keep track in $\mathcal{R}$ of how different cycles are connected. This is not necessary if we reorganise the walks to generate only connected diagrams from the start. Since for generic Hamiltonians we cannot rely on the cancellation of terms with repeated elements, we at least must keep track of the elements visited up to the current step $l$, =$\mathcal{R}=\{e_{1},e_{2},\ldots e_{l}\}$, and ban adding $e$ to $\mathcal{C}$ if $e\in\mathcal{R}$. Demoting the role of $h$ in the node label $[l,h,e]\otimes\{e_{1},e_{2},\ldots e_{l-1},e\}$ to being merely the first element in the current cycle, we can generate only connected diagrams if each new cycle starts with an element that is paired by an interaction to one of the already visited ones $e_{i}\in\{e_{1},e_{2},\ldots e_{l-1},e\}$, e.g. the smallest in $\mathcal{R}$ that is not already paired, for uniqueness. It is easy to see that the number of floating point operations in this graph is only exponential, $\mathcal{O}(n^{3})4^{n}$, and that the information about visited elements carried in $\mathcal{R}$ is minimal for this order of traversing the sequences $\mathcal{C}$, i.e. the graph cannot be pruned any further. Indeed, at a given level $l$ the number of nodes, i.e. unique labels $[l,h,e]\otimes\mathcal{R}$, is a product of the number of values of $h$, $e$ and $\mathcal{R}$, given by $\mathcal{O}(l)$, $\mathcal{O}(2n-l)$ and $\mathcal{O}((2n)!/l!(2n-l)!)$ respectively. Each node is linked to $\mathcal{O}(n)$ nodes at the level $l+1$, and thus the total number of edges is obtained by summing $n\,l(2n-l)(2n)!/l!(2n-l)!$ over all levels $l$, resulting in the exponential scaling. The practical computational cost of this algorithm, labelled CoS-2, is shown for the $SU(N)$ case in Fig. 2b.
@@
In systems where there is no non-trivial factor associated with each fermionic loop, as, e.g., in the $SU(2)$ Hubbard model, or for $N=1$, cancellations between cycle covers with repeated elements can still be utilised to reduce the cost further to $\mathcal{O}(n^{3})3^{n}$. To this end, $\mathcal{R}$ only needs to store, for each interaction line, the number of its vertices—zero, one or two—that have already been visited. Since there is no record of which of the two vertices of an interaction has been visited, both options for the element that starts a new cycle need to be allowed, with the cycle cover that ends up repeating the vertex cancelling out, as in the case of the determinant. The complexity of this algorithm is plotted for the $SU(2)$ case in Fig. 2a.
@@
Finally, sums of skeleton (bold-line) diagrams in arbitrary channels can be straightforwardly generated in our approach. For instance, the computational cost of producing an expansion in terms of the full (interacting) Green’s function $G$ and screened interaction $W$  Hedin (1965) by a simple extension of the CoS-2 algorithm is plotted in Fig. 2b as CoS-GW. The challenge of restricting the series to irreducible diagrams in both channels is met here by supplementing the nodes in the CoS-2 graph with a record $\mathcal{R}$ that keeps track of connectivity—as in the CoS-1 approach of Sec. II.2—when any two propagators in a cycle or two interaction lines between different cycles are cut. Curiously, there is no notable cost increase relative to the CoS-1 algorithm for connected diagrams. The versatility of the CoS platform could enable more efficient algorithms for skeleton series in the future.
@@ II.4 Vector variant and quantum speed-up
Figure 3: Number of floating-point operations (FLOP) required to evaluate the sum of the integrands of all Feynman diagrams of order $n$ by the CoS-2 algorithm without inherent symmetrisation over vertex permutations: for connected diagrams in the $SU(N)$ case with ordered end points of interactions (CoS-2, ordered interaction ends) and fully ordered vertices (CoS-2, all vertices ordered). Shown as CoS-GW, all vertices ordered is the computational cost of summing the nonsymmetrised skeleton (bold-line) series in terms of the renormalised Green’s function $G$ and screened interaction $W$. Source data are provided as a Source Data file.
@@ II.5 Nonsymmetrised sum
Like the determinant, in its general form the CoS approach sums not just all diagram topologies of a given order $n$ but their $2^{n}n!$ realisations with all permutations of the interaction lines and end points. An advantage of the explicit summation is that this symmetrisation can be suppressed if useful with a reduction of the computational cost. Specifically, since the Hubbard interaction is local in space and time, the symmetrisation over $2^{n}$ exchanges of the interaction end points does not affect the variance of the Monte Carlo sampling. In this case, ordering the arguments in each $V(\alpha,\beta)$ leads to a tangible speed-up of the DiagMC sampling. Indeed, by requiring that a node $[l,h,\bar{e}]\otimes\mathcal{R}$ is added to the graph only if $e$ (the vertex connected to $\bar{e}$ by an interaction) is already present in $\mathcal{R}$, the number of operations in the CoS-2 approach drops from $\mathcal{O}(n^{3}4^{n})$ to $\mathcal{O}(n^{3}3^{n})$. The actual computational cost shown in Fig. 3 (CoS-2, ordered interaction ends) follows this scaling.
@@
II.6 Equation of state of the $2D$ $SU(N)$ Hubbard model
The recent study by Pasqualetti et al. Pasqualetti et al. (2024) has revealed a perfect agreement between the DQMC calculations and experimental measurements of the EoS—the average particle number $\langle n\rangle$ per lattice site as a function of the chemical potential $\mu$—of the $2D$ $SU(N)$ Hubbard model at temperatures down to $T/t=0.3$ at $U/t=2.3$ and coupling up to $U/t=10.4$ at $T/t=1.35$ for $N=6$. As a benchmark, we obtain the $\langle n\rangle(\mu)$ curve at $T/t=0.3$, $U/t=2.3$, plotted in Fig. 4, and find it to be in perfect agreement  [^2] The EoS reported in Ref. Pasqualetti et al. (2024) is averaged over small intervals of the chemical potential due to the experimental resolution, but we have verified the agreement within combined error bars for the prime numerical data, kindly provided by the authors. with that computed and measured in Ref. Pasqualetti et al. (2024).
Ibarra-García-Padilla et al. Ibarra-García-Padilla et al. (2021) demonstrate that the sign problem in their DQMC simulations rapidly intensifies with lowering $T$ and increasing $U$ and $N$ at the considered densities, as long as the system remains compressible. To explore the more challenging regime, the EoS was obtained by DiagMC at a lower temperature $T/t=0.15$ (see Figure 4), for which the $\langle n\rangle(\mu)$ curve is below that for $T/t=0.3$ at $U/t=2.3$, indicating that the entropy density $s(\mu)$ is an increasing function in this range of $\mu$ due to the Maxwell relation $\partial s/\partial\mu=\partial n/\partial T$. Following Ref. Lenihan et al. (2021), this also suggests that the system is in the metallic state in this regime. We further evaluate the series at larger values of $U$ up to $U=12$. A plateau around $\langle n\rangle=1$ can be seen to start emerging at $U=8$, consistently with the development of a (pseudo-)gapped state at strong coupling. The plateau appears more pronounced at $U=12$, but at these couplings the systematic error of resummation beyond the convergence radius (see Methods) becomes comparable to the propagated statistical error, and the combined error bar, i.e. a sum of the two errors, shown in Fig. 4 grows substantially. Nonetheless, the analytic structure of the series appears free from singularities near a positive real $U$, such as those in the $SU(2)$ Hubbard model at similar parameters Kim et al. (2020); Lenihan et al. (2021). There, the growth of the antiferromagnetic (AFM) correlation length beyond $\sim 10$ lattice sites was shown to be responsible for a near-critical behaviour of the diagrammatic expansions at these temperatures and $\langle n\rangle=1$ already at $U/t\sim 3$ (see the Supplemental Material of Ref. Lenihan et al. (2021) for details). Also in contrast to the $SU(3)$ case, where an insulating AFM ground state at $\langle n\rangle=1$ emerges for $U/t\gtrsim 5.5$ Feng et al. (2023) and strong AFM correlations (with a transformation upon heating) are observed up to $T/t\sim 0.5$ at $U/t=8$ Ibarra-García-Padilla et al. (2023), for $N=6$ AFM correlations appear weak down to $T/t=0.15$ at this coupling, as implied by the absence of the corresponding singularities or a pronounced plateau in $\langle n\rangle(\mu)$. This is not unexpected since for larger $N$ the Fermi surface at $\langle n\rangle=1$ is farther from being nested.
Figure 4: Equation of state for the $2D$ $SU(N)$ Hubbard model with $N=6$ for $T/t=0.3,0.15$ and $U/t=2.3,4,8,12$. The error bars are a sum of the systematic error of resummation beyond the convergence radius and the propagated statistical (standard deviation) error of the series coefficients; see Methods. Source data are provided as a Source Data file.
III Discussion
The example application to the EoS of the $2D$ $SU(6)$ Hubbard model provides controlled benchmarks for ongoing theoretical and experimental studies, aimed at accessing lower temperatures and novel quantum many-body states. As a byproduct of a diagrammatic calculation, the analytic structure of the series offers additional insights in the physics of the system. The results suggest a phase transition in the attractive $SU(N)$ Hubbard model at a coupling strength as low as $U_{c}/t\sim-1$ up to temperatures $T/t\lesssim 0.5$ (see Methods), and absence of evidence for strong AFM correlations in the repulsive case at the considered temperatures and interaction strengths, at which the $SU(2)$imkovic et al. (2020); Kim et al. (2020) and $SU(3)$ Ibarra-García-Padilla et al. (2023); Feng et al. (2023) Hubbard models are already in the (quasi-)AFM state. In the $SU(2)$ case, the formulation in the thermodynamic limit enabled DiagMC to attain controlled accuracy in the regime where correlations are intrinsically long-range and are difficult to capture reliably by finite-size methods even in the absence of the sign problem Šimkovic et al. (2020); Kim et al. (2020). Such regimes of the $SU(N)$ model is where the developed technique can prove particularly useful. The possibility of a direct calculation of entropy in the DiagMC approach Lenihan et al. (2021) could be instrumental for thermometry in experiments with ultracold atoms that are currently testing the limits of state-of-the-art theoretical methods.
@@ III Discussion
IV Methods
The average particle number per lattice site $\langle n\rangle$ is expressed in our approach as an expansion in the powers of the Hubbard coupling $U$, $\langle n\rangle(T,\mu,U)=\sum_{m=0}^{\infty}a_{m}(T,\mu)U^{m}$, in the thermodynamic limit. In DiagMC $\mu$ is typically shifted to improve convergence Wu et al. (2017); a $U$-dependent shift, described, e.g., in Ref. Šimkovic and Kozik (2019) is adopted here. The series coefficients $a_{m}$ are obtained by the Monte Carlo integration of all connected diagram topologies of order $m$—summed here by the CoS-2 technique with ordered interaction ends in $\mathcal{O}(m^{3}3^{m})$ operations—over the positions of the $m$ interaction lines in space-imaginary-time. Thus, $a_{m}$ are known numerically exactly with statistical error bars, while the only source of systematic error is the truncation of the series at order $n$. Although the series turns out to be divergent in all regimes of interest, being able to evaluate $a_{m}$ up to $n=8$ with $<5\%$ statistical error (and fractions of percent at lower orders) enables an accurate reconstruction of the result with controlled precision.
@@ IV Methods
Fig. 5a shows the partial sum for the density of the $2D$ $SU(6)$ Hubbard model at $U/t=2.3$, $\mu/t=0.575$ and $T/t=0.3$, the lowest temperature in the study by Pasqualetti et al. Pasqualetti et al. (2024), as a function of the truncation order $n$. The series is seen to wildly diverge, but its analytic structure is rather simple, with the dominating singularity at $U_{c}/t\approx-0.9(1)$, which allows us to reconstruct the result following the approach developed in Ref. Šimkovic and Kozik (2019). Specifically, we employ the Dlog-Padé Baker Jr (1961) technique, taking into account the statistical error bars of $a_{m}$ and verifying that the systematic error of the evaluation—detected by the variation of the answer with free parameters of Dlog-Padé—is negligible in this regime. The result for the series in Fig. 5 is $\langle n\rangle=1.335(5)$. The singularity at a real $U$ is indicative of a phase transition exhibited by the attractive $SU(N)$ Hubbard model, which is likely to a superfluid state. When the series for the relevant susceptibility is considered, the divergence at $U_{c}$ is an accurate tool for characterising the critical point Lenihan et al. (2022). Leaving the calculation of susceptibilities for a more focused study, we plot in Fig. 5b a crude estimate of $U_{c}$ from the divergence of density at $T/t=0.3$.
@@ Acknowledgements.
The author is grateful to Kaden Hazzard for illuminating and stimulating discussions, to Eduardo Ibarra-García-Padilla, Sohail Dasgupta, Kaden Hazzard, and Richard Scalettar for sharing their DQMC data, and to Sohail Dasgupta, Simon Fölling, Kaden Hazzard, Eduardo Ibarra-García-Padilla, Giulio Pasqualetti, and Richard Scalettar for a fruitful exchange of results and ideas. This work was supported by EPSRC through Grant No. EP/X01245X/1. The calculations were performed using King’s Computational Research, Engineering and Technology Environment (CREATE).